The following information is provided to assist you in evaluating the performance of the production operations of Studio Company:

Units produced (actual) 55,000
Master production budget
Direct materials $128,040
Direct labor 108,640
Overhead 178,480
Standard costs per unit
Direct materials $1.65 × 2 gallons per unit of output
Direct labor $14 per hour × 0.2 hour per unit
Variable overhead $13.00 per direct labor-hour
Actual costs
Direct materials purchased and used $141,050 (80,600 gallons)
Direct labor 133,497 (9,780 hours)
Overhead 175,200 (61% is variable)

Variable overhead is applied on the basis of direct labor-hours.

Required:

Calculate all variable production cost price and efficiency variances and fixed production cost price and production volume variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

Price Variance Efficiency Variance Production Volume Variance
Direct materials
Direct labor
Variable overhead
Fixed overhead
Answer #1

Working notes:

Budgeted units = $128,040 / 1.65*2 = 38,800 units

Direct material price variance = Actual quantity * standard price - Actual quantity * Actual price

Direct material price variance = 80,600*$1.65 - 141,050 = $8,060 Unfavorable

Direct material efficiency variance = Standard quantity*Standard price - Actual quantity * standard price

Direct material efficiency variance = 55,000*2*$1.65 - 80,600*$1.65 = $48,510 Favorable

Direct labor price variance = Actual hours * standard price - Actual hours * Actual price

Direct labor price variance = 9,780*$14 - $133,497 = $3,423 Favorable

Direct labor efficiency variance = Standard hours*Standard price - Actual hours * standard price

Direct labor efficiency variance = 55,000*0.2*$14 - 9,780*$14 = $17,080 Favorable

Variable overhead price variance = Actual hours * standard price - Actual hours * Actual price

Variable overhead price variance = 9,780*$13 - $106,872 (175,200*61%) = $20,268 Favorable

Variable overhead efficiency variance = Standard hours*Standard price - Actual hours * standard price

Variable overhead efficiency variance = 55,000*0.2*$13 - 9,780*$13 = $15,860 Favorable

Fixed overhead price variance = Actual overhead - Budgeted overhead

Budgeted overhead = $178,480 - 100,880 (38,800*0.2*$13) = $77,600

Fixed overhead price variance = $68,328 (175,200*39%) - 77,600 = $9,272 Favorable

Fixed overhead production volume variance = Budgeted overhead - Absorbed overhead

Absorbed overhead = $77,600 / 38,800 * 55,000 = $110,000

Fixed overhead production volume variance = $77,600 - 110,000 = $32,400 Favorable

Price variance Efficiency variance Production volume variance
Direct materials $8,060 U $48,510 F
Direct labor 3,423 F 17,080 F
Variable overhead 20,268 F 15,860 F
Fixed overhead 9,272 F $32,400 F